please help me answer this problem: 3 years ago, Dina was 1/3 of her mother's age.In 7 years, she will be 1/2 of her mother's age. How old is Dina now?
Answer by mananth(16946)

3 years ago, Dina was 1/3 of her mother's age.In 7 years, she will be 1/2 of her mother's age. How old is Dina now?
...
mother's age be 3x
Dina's age =x
..
3 years ago
mother's age = 3x-3
Dina's age = x-3
..
after 7 years
mother's age = 3x-3+10
3x+7
Dina's age =x-3+10
x+7
x+7=1/2*(3x+7)
2(x+7)=3x+7
2x+14=3x+7
x=7 Dina's age
...
m.ananth@hotmail.ca